OpenStudy (anonymous):

How do I solve: what is 15% of 80...Explaining the steps would help.

OpenStudy (michele_laino):

hint: you have to do this computation please: \[\huge \frac{{15}}{{100}} \times 80 = ...?\] since the exercise, asks for a portion of those 80 units

OpenStudy (anonymous):

I get different answers...over and above the number eighty.

OpenStudy (michele_laino):

what are your options, please?

OpenStudy (anonymous):

There are no options

OpenStudy (aaronandyson):

15% of 80 : \[15 \times 80 \over 100\]

OpenStudy (aaronandyson):

15*8 = 120 So whats 15*80..?

OpenStudy (anonymous):

15*80 = 960

OpenStudy (aaronandyson):

no..! Please show work our work..

OpenStudy (aaronandyson):

your*

OpenStudy (aaronandyson):

Here's an example: 5 time 5 is 25 5 times 50 is 250 5 times 500 is 2,500 5 times 5000 is 25,000 Yous seeing the trend..?

OpenStudy (anonymous):

I see the trend.

OpenStudy (aaronandyson):

So 15 times 8 is 120 Using the example above whats 15 times 80..?

OpenStudy (anonymous):



OpenStudy (anonymous):

I tried to use the drawing...not as neat as I would like it to be.

OpenStudy (aaronandyson):

No, no,no..! Read the example above once again.

OpenStudy (aaronandyson):

15*8 is 120 15*80 is ?

OpenStudy (anonymous):

1200

OpenStudy (aaronandyson):

CORRECT!

OpenStudy (aaronandyson):

Okay . Moving on: \[15 \times 80 \over 100\] = \[1200 \over 100\]

OpenStudy (aaronandyson):

cancel the zeros.

OpenStudy (anonymous):

\[\frac{ 12 }{ 1}\]

OpenStudy (aaronandyson):

CORRECT!

OpenStudy (aaronandyson):

\[12/1\] biols down to just 12

OpenStudy (aaronandyson):

boils*

OpenStudy (anonymous):

Cool, thanks.

OpenStudy (aaronandyson):

Your welcome :D

OpenStudy (shadowlegendx):

*youʻre